OpenStudy (litzyceron):

Which equation is represented by the table? x y (x, y) 0 3 (0, 3) 1 4 (1, 4) 2 5 (2, 5) A. y = 5x +3 B. y = 3x – 5 C. y = x + 3 D. y = 2x + 2

random231 (random231):

again put in the values and see which equation it satisfies.

OpenStudy (litzyceron):

???

OpenStudy (litzyceron):

@mikey35633

OpenStudy (anonymous):

whats up

OpenStudy (litzyceron):

Which equation is represented by the table? x y (x, y) 0 3 (0, 3) 1 4 (1, 4) 2 5 (2, 5) A. y = 5x +3 B. y = 3x – 5 C. y = x + 3 D. y = 2x + 2

OpenStudy (anonymous):

ok hold on i got it

OpenStudy (litzyceron):

ok

OpenStudy (anonymous):

The answer is C. (0, 3) y = x + 3 3 = 0 + 3 3 = 3 (1, 4) y = x + 3 4 = 1 + 3 4 = 4 (2, 5) y = x + 3 5 = 2 + 3 5 = 5

OpenStudy (anonymous):

option c

OpenStudy (anonymous):

what next

OpenStudy (litzyceron):

ok hold on